Chi-Square Data

Brown Eyes

Bushy Eyebrows

Brown Eyes

Thin Eyebrows

Blue Eyes

Bushy Eyebrows

Blue Eyes

Thin Eyebrows

Observed

5

7

0

0

Expected

4.5

4.5

1.5

1.5

Show your Chi-Square Calculations below:

The p value from the analysis is _______.  The degrees of freedom are __________.

The results show (check on circle):

The Observed and Expected are significantly different

The Observed and Expected are not significantly different

Discuss what your results tell us about how these two genes are placed on the same chromosome.
Answer #1

Degrees of freedom formula = n-1

Where is n is the number of classes.

Therefore df= 4-1=3 .

\chi 2=(Observed-expected)2/expected value.

= (5-4.5)2/4.5+(7-4.5)2/4.5+(0-1.5)2/1.5+(0-1.5)2/1.5

= 0.05+1.38+1.5+1.5=4.43

In a chi square table the p value is in between 0.25 and 0.10 at a level of significance is 0.05. since the p value is greater than the significance value then the null hypothesis is accepted.

By accepting the null hypothesis we can conclude that the two genes are placed on the same chromosome.